### v1

# Quantum dissipative dynamics of driven Duffing oscillator near attractors

Abstract
We investigate the quantum dissipative dynamics near the stable states (attractors) of a driven Duffing oscillator. A refined perturbation theory that can treat two perturbative parameters with different orders is developed to calculate the quantum properties of Duffing oscillator near the attractors. We obtain the perturbative analytical results for the renormalized level spacings and the quantum-fluctuation-induced distribution of higher energy levels near the attractor, which are then verified by numerical simulations. Furthermore, we demonstrate that strong damping induces a slight renormalization of level spacing, and the renormalized Bose distribution affected by dephasing. Our work provides new insights into the quantum dynamics of driven Duffing oscillator and offers a theoretical framework that can be applied to related quantum system near their stable states.
I Introduction
The driven Duffing oscillator Duf18 , a paradigmatic model for various nonlinear mechanical systems and nonlinear optical phenomena, has been fascinating physicists for a long time by its rich dynamical behaviors such as bistability, bifurcation, and chaotic trajectories. In recent years, the Duffing oscillator has received renewed attention as the quantum regime of nanomechanical oscillators becomes experimentally accessible. The interplay between quantum effects, nonequilibrium dynamics, and nonlinear effects makes the driven Duffing oscillator a model with broad applicability across multiple domains, e.g., mechanical metrology Dyk12m ; Cle02 ; Lif08 ; Poo12 , cavity and circuit quantum electrodynamics Jea24 ; Mav16 ; Boi10 , nano- and opto-mechanics Buk06 ; Alm07 ; Del07 ; Kip11 ; Dyk22 ; Pet14 ; Bol23 ; Hel24 ; Dyk22R , and cold atoms Art03 ; Got19 ; Par12 . Notable examples include bifurcation-based quantum measurement devices, where the low- and high-amplitude states of bistability are entangled with the ground and excited states of qubits respectively, enabling the analysis of the qubit’s state through the detection of classical signals Ald05 ; Sid04 ; Sid06 ; Man07 ; Sid09 .
The dissipative dynamics of driven Duffing oscillators have been extensively studied. In the underdamped regime near a bifurcation point, a scaling law for the noise-induced escape from metastable states was established Dyk05 . It was also revealed that the quantum activation process has distinct temperature dependency compared to that for the quantum tunneling process Dyk07 . The distinct transition rate scaling behaviors near bifurcation points were also revealed in the driven mesoscopic Duffing oscillator Guo10 . It was found that the bifurcation point is shifted by the quantum effect and a linear scaling behavior for the tunneling rate with the driving distance to the shifted bifurcation point Guo11 . Recent advances also showed that the quadrature squeezing can enhance the Wigner negativity in a Duffing oscillator, demonstrating a promising approach to generate nonclassical states in macroscopic mechanical systems Ros24 . For two Duffing oscillators coupled via nonlinear interactions, the stationary paired solutions and their dynamical stability were demonstrated Hel24 . In a coupled system consisting of a time-delayed Duffing oscillator (as a driver system) and a non-delayed Duffing oscillator (as a response system), the phenomenon of transmitted resonance was investigated Coc24 .
In this paper, we focus on investigating quantum dissipative dynamics near the attractors of a driven Duffing oscillator. We develop an effective quantum master equation that can address quantum fluctuations, thermal effects, damping, and dephasing in a unified framework. We quantify the occupation of high levels near the bottom of the potential well by the the quantum squeezing effects. We also demonstrate the effects of strong damping and dephasing on the system’s dynamics, including level spacing renormalization and dephasing-modified Bose distributions. By comparing our theoretical predictions with exact numerical simulations, we demonstrate the accuracy and utility of our proposed framework.
II General Theory
Figure 1: (a) Quasienergy landscape in phase space of a driven Duffing oscillator. The extrema correspond to the high amplitude stable state (M) and the low amplitude stable state (m), while the saddle point (s) denotes an unstable state. (b) Cross-section of the quasienergy potential at $P=0$. Quantum energy levels close to the maximum (M) and minimum (m) of the potential are depicted by blue and red lines, respectively. The unstable saddle point (s) is also indicated.
II.1 Model Hamiltonian
An extensive class of macroscopic physical systems, such as Josephson junctions and nanomechanical oscillators can be modeled by the Duffing oscillator in the presence of a periodic driving force with the system Hamiltonian described by Ser07
${H}_{S}(t)=\frac{{p}^{2}}{2m}+\frac{1}{2}m\Omega^{2}{x}^{2}-\gamma{x}^{4}+F(t){x}.$ (1)
Here, parameter $m$ ($\Omega$) describes the mass (frequency) of the oscillator, $\gamma$ gives the nonlinearity of Duffing oscillator, and $F(t)=F_{0}(e^{i\nu t}+e^{-i\nu t})$ describes the periodically driving force with frequency $\nu$. By switching to the rotating frame using the transformation $U(t)=\exp(-i\nu a^{\dagger}at)$, where $a^{\dagger}$ and $a$ are the oscillator raising and lowering operators, and applying the rotating wave approximation (RWA), we obtain a time-independent Hamiltonian
$H=(\delta\omega+\chi)a^{\dagger}a+\chi(a^{\dagger}a)^{2}+\epsilon(a^{\dagger}+a)\,,$ (2)
where $\delta\omega=\hbar(\Omega-\nu)$ is the frequency detuning, $\chi=-3\gamma\hbar^{2}/2(m\Omega)^{2}$ is the scaled dimensionless nonlinearity, and $\epsilon=F_{0}\sqrt{{\hbar}/{2m\Omega}}$ is the scaled driving strength. We then introduce the position operator $Q$ and momentum operator $P$ in the rotating frame via
$Q=\sqrt{\frac{\lambda}{2}}(a^{\dagger}+a),\ \ \ P=i\sqrt{\frac{\lambda}{2}}(a^{\dagger}-a),$ (3)
which satisfy the commutation relation
$[Q,P]=i\lambda.$ (4)
Here, the parameter $\lambda=-\chi/(4\Delta)$ is the dimensionless Planck constant that describes the quantumness of the system, i.e., the value of $\lambda$ increases as the system approaches the quantum regime. Substituting operators $Q$ and $P$ back into the RWA Hamiltonian (2), we obtain
$H/\Delta=\frac{1}{\lambda}(g+\frac{1}{4})-\frac{1}{2}+\frac{\lambda}{4},$ (5)
where $g$ is the quasienergy given by
$g=-(Q^{2}+P^{2}-1)^{2}/4+\sqrt{\beta}Q.$ (6)
Here, the parameter $\beta=-f^{2}\chi/(2\Delta^{3})$ is the scaled driving strength. Note that Eq. (6) is valid only for the soft nonlinearity $\chi>0$. For the hard nonlinearity $\chi<0$, the quasienergy is given by $g=(Q^{2}+P^{2}-1)^{2}/4-\sqrt{\beta}Q$.
II.2 Renormalized master equation
The characteristic behavior of the driven Duffing system is the bistability manifesting as two stable states: the low-amplitude state (LAS) and the high-amplitude state (HAS). As depicted in Fig. 1(a), these stable states correspond to the extrema in the quasienergy landscape, which are defined as the attractors in the phase space. The unstable state, known as the saddle point, is located on the separatrix, serving as the boundary dividing the basins of the attractors. When damping is present, the system evolves towards the nearby attractor if its initial state lies within the basin of the attractor. However, due to thermal noise, the system does not remain exactly on the attractor, but forms a probability distribution within its basin of attraction. To describe the dissipative dynamics of our system, we employ a Lindblad form master equation
$\frac{d\rho}{dt}=-i[H,{\rho}]+\frac{\kappa}{2}\{(1+\bar{n}){\cal D}[a]{\rho}+\bar{n}{\cal D}[a^{\dagger}]{\rho}\},$ (7)
where ${\cal D}[\bullet]$ is the Lindblad operator defined as ${\cal D}[A]{\rho}\equiv 2A{\rho}A^{\dagger}-A^{\dagger}A{\rho}-{\rho}A^{\dagger}A$, $\bar{n}$ is the Bose-Einstein distribution, and $\kappa$ is the damping strength.
To study the quantum dynamics near the bottom of each stable state, we first transform the system to the center of the attractor using the displacement operator
$D[\alpha]=e^{\alpha a^{\dagger}-\alpha^{*}a}$ (8)
with parameter $\alpha$ a complex number. By defining the displaced density matrix $\tilde{\rho}=D[\alpha]^{\dagger}\rho D[\alpha]$, we obtain the following master equation
$\frac{d\tilde{\rho}}{dt}$ $=$ $-i\left[\tilde{H},{\tilde{\rho}}\right]+\frac{\kappa}{2}\{(1+\bar{n}){\cal D}[a]{\tilde{\rho}}+\bar{n}{\cal D}[a^{\dagger}]{\tilde{\rho}}\}$ (9)
$+\left[{\tilde{\alpha}}^{*}a-{\tilde{\alpha}}a^{\dagger},\tilde{\rho}\right]\,,$
with ${\tilde{\alpha}}=[\frac{1}{2}\kappa+i(\delta\omega+\chi+2\chi\alpha|^{2})]\alpha+i\varepsilon.$ By choosing $\alpha$ such that $\tilde{\alpha}=0$, the master equation is simplified into
$\frac{d\tilde{\rho}}{dt}$ $=$ $-[\tilde{H},{\tilde{\rho}}]+\frac{\kappa}{2}\{(1+\bar{n}){\cal D}[a]{\tilde{\rho}}+\bar{n}{\cal D}[a^{\dagger}]{\tilde{\rho}}\}$ (10)
with the renormalized Hamiltonian $\tilde{H}$ given by
$\tilde{H}$ $=$ $(\delta\omega+4\chi|\alpha|^{2})a^{\dagger}a+\chi(a^{\dagger}a)^{2}+\chi({\alpha^{*}}^{2}a^{2}+{\alpha}^{2}{a^{\dagger}}^{2})$ (11)
$+2\chi(\alpha{a^{\dagger}}^{2}a+\alpha^{*}{a^{\dagger}}a^{2}).$
We then introduce the squeezing operator
$S(\xi)=e^{(\xi^{*}a^{2}-\xi{a^{\dagger}}^{2})/2},$ (12)
which has the transformation property $S^{\dagger}aS=va+ua^{\dagger}$ with $v=\cosh(|\xi|)$ and $u=-\frac{|\xi|}{\xi}\sinh(|\xi|)$. Bu defining the squeezed density operator $\bar{\rho}=S^{\dagger}\tilde{\rho}S$, we transform the master equation Eq. (10) into the following form
$\frac{d\bar{\rho}}{dt}$ $=$ $-i[\bar{H},\bar{\rho}]+\frac{\kappa}{2}\{(1+\bar{N}){\cal D}[a]\bar{\rho}+\bar{N}{\cal D}[a^{\dagger}]\bar{\rho}\}$ (13)
$+\frac{\kappa}{2}M(2a^{\dagger}\bar{\rho}a^{\dagger}-{a^{\dagger}}^{2}\bar{\rho}-\bar{\rho}{a^{\dagger}}^{2})$
$+\frac{\kappa}{2}M^{*}(2a\bar{\rho}a-{a}^{2}\bar{\rho}-\bar{\rho}{a}^{2})$
$-i[\bar{\xi}{a^{\dagger}}^{2}+\bar{\xi}^{*}{a}^{2},\bar{\rho}],$
where $\bar{N}=\bar{n}|v|^{2}+(1+\bar{n})|u|^{2}$ is the effective Bose distribution and $M=uv^{*}(2\bar{n}+1)$ is the squeezing number. The parameter $\bar{\xi}$ in the last term of Eq. (13) is given by
$\bar{\xi}=[\delta\omega+4\chi|\alpha|^{2}+2\chi(2|u|^{2}+|v|^{2})]v^{*}u+\chi({\alpha^{*}}^{2}u^{2}+\alpha^{2}{v^{*}}^{2}).$
By setting $\bar{\xi}=0$, the renormalized master equation (13) is further simplified into
$\frac{d\bar{\rho}}{dt}$ $=$ $-i[\bar{H},\bar{\rho}]+\frac{\kappa}{2}\{(1+\bar{N}){\cal D}[a]\bar{\rho}+\bar{N}{\cal D}[a^{\dagger}]\bar{\rho}\}$ (14)
$+\frac{\kappa}{2}M(2a^{\dagger}\bar{\rho}a^{\dagger}-{a^{\dagger}}^{2}\bar{\rho}-\bar{\rho}{a^{\dagger}}^{2})$
$+\frac{\kappa}{2}M^{*}(2a\bar{\rho}a-{a}^{2}\bar{\rho}-\bar{\rho}{a}^{2})\,.$
The final renormalized Hamiltonian $\bar{H}$ becomes
$\bar{H}$ $=$ $\delta\bar{\omega}a^{\dagger}a+\bar{\chi}(a^{\dagger}a)^{2}$
$+2\chi S^{\dagger}(\alpha{a^{\dagger}}^{2}a+\alpha^{*}{a^{\dagger}}a^{2})S+\chi F,$
with $F=2(|v|^{2}+|u|^{2})(v^{*}u{a^{\dagger}}^{3}a+vu^{*}a^{\dagger}a^{3})+(v^{*}u)^{2}{a^{\dagger}}^{4}+(u^{*}v)^{2}{a}^{4}$. The renormalized detuning $\delta\bar{\omega}$ and nonlinearity $\bar{\chi}$ are given by
$\left\{\begin{array}[]{lll}\delta\bar{\omega}/\delta\omega&=&(1-4\lambda|\alpha|^{2})(|v|^{2}+|u|^{2})\\ &&-2\lambda(2|uv|^{2}+|u|^{4}+{\alpha^{*}}^{2}uv+\alpha^{2}u^{*}v^{*}),\\ \bar{\chi}/\delta\omega&=&-\lambda(|u|^{4}+|v|^{4}+4|uv|^{2})\end{array}\right.$ (19)
with the displacement parameter $\alpha$, the squeezing parameters $u$ and $v$ satisfying the following steady equations
$\left\{\begin{array}[]{lll}&0=&[\frac{\kappa}{2\delta\omega}+i(1-\lambda-2\lambda|\alpha|^{2})]\alpha-i\sqrt{\frac{\beta}{2\lambda}},\\ &0=&[1-4\lambda|\alpha|^{2}-\lambda(4|u|^{2}+2|v|^{2})]v^{*}u\\ &&-\lambda({\alpha^{*}}^{2}u^{2}+\alpha^{2}{v^{*}}^{2}).\end{array}\right.$ (23)
Here, we have introduced the dimensionless driving strength $\beta=2\lambda(\varepsilon/\delta\omega)^{2}$.
II.3 Orders of perturbative parameters
To perform perturbation calculations for the renormalized Hamiltonian (II.2), we can choose dimensionless Planck constant $\lambda=-\chi/(4\Delta)$ as the natural choice for the perturbation parameter. However, since the displacement parameter $\alpha$ is also a function of $\lambda$, it is subtle to properly organize the perturbative terms according to their respective orders. In fact, the Hamiltonian (II.2) should be written in different forms for different attractors. The stable state of the driven Duffing oscillator can be approximated as a coherent state $|\alpha\rangle$. By applying the variational principle in quantum mechanics, i.e., $\partial_{\alpha}\langle\alpha|H|\alpha\rangle=0$, we obtain two solutions of the steady coherent number: a smaller one $|\alpha_{l}|^{2}\approx\beta/(2\lambda)$ for the LAS and a larger one $|\alpha_{h}|^{2}\approx 1/(2\lambda)$ for the HAS.
For the LAS, we sort the terms in the renormalized Hamiltonian (II.2) as follows
$\frac{\bar{H}^{l}}{\delta\omega}=h^{l}_{0}+\lambda h^{l}_{\lambda}+\beta h^{l}_{\beta}+\sqrt{\lambda\beta}h^{l}_{\lambda\beta},$ (24)
with each term given by
$h^{l}_{0}$ $=$ $(|v|^{2}+|u|^{2})a^{\dagger}a,$
$h^{l}_{\lambda}$ $=$ $-2(2|uv|^{2}+|u|^{4})a^{\dagger}a$
$-(|u|^{4}+|v|^{4}+4|uv|^{2})(a^{\dagger}a)^{2}-F,$
$h^{l}_{\beta}$ $=$ $-\frac{1}{\beta}[{4\lambda|\alpha_{l}|^{2}}(|v|^{2}+|u|^{2})$
$+\lambda(2{\alpha^{*}_{l}}^{2}uv+2\alpha^{2}_{l}u^{*}v^{*})]a^{\dagger}a,$
$h^{l}_{\lambda\beta}$ $=$ $-2\sqrt{{\lambda}/{\beta}}\ S^{\dagger}(\alpha_{l}{a^{\dagger}}^{2}a+\alpha^{*}_{l}{a^{\dagger}}a^{2})S\,.$ (25)
Together with Eq. (23), we can now perform perturbative calculations near the bottom of the LAS. Given that the dimensionless driving strength $\beta$ is also small, we consider the sum of the last three terms in the Hamiltonian Eq. (24) as the perturbation term and calculate the desired quantities perturbatively.
For the HAS, as the coherent number $|\alpha_{h}|^{2}\approx 1/(2\lambda)$ can be significantly large for $\lambda\ll 1$, a more careful sorting of the terms in the renormalized Hamiltonian Eq. (II.2) is needed, along with the steady-state condition Eq. (23), to ensure terms of the same order are kept together. We introduce $\gamma=\sqrt{\lambda}$ as the perturbation parameter and rewrite the Hamiltonian as
$\frac{\bar{H}^{h}}{\delta\omega}$ $=$ $h^{h}_{0}+\gamma h^{h}_{1}+\gamma h^{h}_{2}$ (26)
with each term given by
$h^{h}_{0}$ $=$ $[(1-4\lambda|\alpha_{h}|^{2})(|v|^{2}+|u|^{2})$
$-2\lambda({\alpha^{*}_{h}}^{2}uv+\alpha_{h}^{2}u^{*}v^{*})]a^{\dagger}a\,,$
$h^{h}_{1}$ $=$ $-\sqrt{{\lambda}}\ S^{\dagger}(2\alpha_{h}{a^{\dagger}}^{2}a+2\alpha^{*}_{h}{a^{\dagger}}a^{2}+\alpha_{h}{a^{\dagger}}+\alpha^{*}_{h}a)S\,,$
$h^{h}_{2}$ $=$ $-2(2|uv|^{2}+|u|^{4})a^{\dagger}a-(|u|^{4}+|v|^{4}+4|uv|^{2})(a^{\dagger}a)^{2}$ (27)
$-F-2(2|u|^{2}+|v|^{2})(v^{*}u{a^{\dagger}}^{2}+u^{*}v{a}^{2})\,.$
To handle the perturbation orders coherently, we have rearranged the perturbation terms by removing those of order $\lambda$ from the steady condition Eq. (23) and incorporating them into the renormalized Hamiltonian. The coherent number $\alpha_{h}$, $u$ and $v$ for the HAS are now determined by the revised steady condition
$\left\{\begin{array}[]{lll}&&[\frac{\kappa}{2\delta\omega}+i(1-2\lambda|\alpha_{h}|^{2})]\alpha_{h}-i\sqrt{\frac{\beta}{2\lambda}}=0,\\ &&(1-4\lambda|\alpha_{h}|^{2})v^{*}u-\lambda({\alpha^{*}_{h}}^{2}u^{2}+\alpha_{h}^{2}{v^{*}}^{2})=0.\end{array}\right.$ (29)
The behavior near the bottom of the LAS is relatively simple and can be modeled using a harmonic oscillator. However, for the HAS, the nonlinear term $\chi(a^{\dagger}a)^{2}\approx\chi|\alpha_{h}|^{2}$ becomes prominent, and the oscillator behaves as a highly squeezed coherent state. In the following sections, we will apply our perturbative method to calculate the crucial quantities related to the HAS of the driven Duffing oscillator, namely the level spacing, averaged position of energy level, and the effective temperature in the vicinity of the HAS attractor.
III Results
III.1 Quantum dynamics of HAS
The quantum dynamics of the driven Duffing oscillator near the HAS attractor exhibit a rich interplay between nonlinearity, quantum fluctuations, and thermal noise. In this section, we discuss the quantum properties of the HAS using the renormalized master equation combined with a refined perturbation theory.
III.1.1 Level Spacing
Figure 2: Comparison of perturbation calculations and exact results for energy level spacing. The exact results (lowest curve) for level spacing $\Delta E_{n}=|E_{n+1}-E_{n}|$ are compared with the zeroth-order (constant spacing), the second-order (second highest line, $\gamma^{2}=\lambda$) and the fourth-order (third highest line, $\gamma^{4}=\lambda^{2}$) corrections. Excellent agreement is observed for low level numbers near the potential well bottom, while higher-order corrections are required for large $n$. Parameters: $\lambda=0.016$ and $\beta=4/75$.
The nonlinearity term $\chi(a^{\dagger}a)^{2}$ in the Hamiltonian, having an opposite sign to $\delta\omega$, results in a decrease in the level spacing $\Delta E_{n}=|E_{n+1}-E_{n}|$ as we approach the saddle point, as illustrated in Fig. 1(b). One can calculate the level spacings with standard perturbation theory by treating the sum of $\gamma h_{1}^{h}$ and $\gamma h_{2}^{h}$ in Eq. (26) as one perturbation term. However, it becomes a challenge to control the accuracy of the level spacings by the perturbative parameter. We find it necessary to distinguish between these two perturbative terms in the perturbation calculations to accurately determine the level spacing. To address this, we have developed a double perturbation theory framework that is particularly suited for the HAS Hamiltonian containing second-order small terms, see the details in Appendix A,
In Fig. 2, we compare our perturbation calculations with the exact numerical results obtained by diagonalizing the original Hamiltonian Eq. (2), which shows an excellent agreement for energy levels near the bottom of the potential well. Under the zeroth-order perturbation approximation, the energy level spacing remains constant across all levels, similar to that of harmonic oscillators. The second-order and fourth-order corrections correspond to accuracies up to $\gamma^{2}=\lambda$ and $\gamma^{4}=\lambda^{2}$, respectively. Higher-order perturbation calculations become necessary for levels farther from the bottom.
Figure 3: Average position of energy levels. Average position $\langle N|a|N\rangle$ of the energy level $N$, illustrating the shifts due to quantum fluctuations. Under the harmonic approximation, $\langle N|a|N\rangle$ is a constant for every level (lowest line). Considering higher-order corrections, we observe changes in $\langle N|a|N\rangle$ with levels. Comparing with the second order perturbation result (highest line), the third order result (second highest line) already align well with the exact one (the bold line).
III.1.2 Averaged level position
We denote the eigenstate of the renormalized Hamiltonian $\bar{H}$ as $|n^{\prime}\rangle$ that is generally a superposition of harmonic oscillator eigenstates $|n^{\prime}\rangle=|n\rangle+\sum_{k\neq n}\xi_{kn}|k\rangle$, where the superposition coefficients $\xi_{kn}$ are provided in the Appendix A. The eigenstate $|N\rangle$ of the original Hamiltonian (2) is related to that of the renormalized Hamiltonian (II.2) via the relationship $|N\rangle=DS|n^{\prime}\rangle$. The matrix element $\langle N|a|M\rangle$ for different levels $|N\rangle$ and $|M\rangle$ is then given by
$\langle N|a|M\rangle$ $=$ $\langle n^{\prime}|S^{\dagger}D^{\dagger}aDS|m^{\prime}\rangle$
$=$ $\langle n^{\prime}|(va+u^{\dagger}+\alpha)|m^{\prime}\rangle$
$=$ $v\langle n^{\prime}|a|m^{\prime}\rangle+u\langle m^{\prime}|a|n^{\prime}\rangle^{*}+\alpha\langle n^{\prime}|m^{\prime}\rangle\,.$
The matrix element $\langle N|a|N\rangle$ provides insight into the average position of each level in the phase space. Under the harmonic approximation ($\xi_{kn}=0$), $\langle N|a|N\rangle$ is a constant $\alpha$ for all levels. However, considering higher-order corrections, the average position $\langle N|a|N\rangle$ changes with energy level, as depicted in Fig. 3. The perturbation calculations agree well with numerical results.
Figure 4: Comparison of stationary probability distributions. (a) Comparison of the stationary probability distribution obtained through perturbation theory to zeroth order (red line), second order (blue lines) and exact numerical simulations (black line). (b) Relative error $\Delta p_{n}/p_{n}$ in the stationary distribution obtained from perturbation theory compared to numerical results. The discrepancy increases for higher levels but remains less than 1.
Figure 5: Ratio of probabilities for the lowest levels in (a) the LAS and (b) HAS, plotted as a function of the Bose distribution $\bar{n}$. The agreement between the analytical results (black square dots) and the exact numerical results (red circle dots) is excellent.
Figure 6: Effective temperature of energy levels. The stationary probability distributions $N_{eff}(n)$ of level $|n\rangle$ obtained through perturbation theory to zeroth order (red line) and second order (blue lines) are compared with the exact numerical simulations (black line). The zero-order effective temperature term gives a constant effective temperature. The second order results in changes in $N_{eff}(n)$, which is quite accurate for levels near the bottom.
III.1.3 Effective temperature
Next, we calculate the stationary distribution over the levels of the HAS and the effective temperature near the bottom. It is important to note that the annihilation operator $a$ is for the Fock state $|N\rangle$, which decreases the Fock state from a higher level to the next lower level $a|N\rangle=\sqrt{N}|N-1\rangle$. However, in our case, the eigenstate of quasienergy $|n\rangle$ is the superposition of the Fock state $|N\rangle$. As a result, the annihilation operator $a$ can either decrease or increase the state $|n\rangle$ even at zero temperature.
Under the assumption of weak damping ($\kappa\ll E_{n}-E_{n+1}$), the off-diagonal matrix elements on the state $|n\rangle$ are very small. Thus, we can only keep the diagonal elements. Here, we assume that the stationary density matrix is diagonal and denote the diagonal terms $p_{n^{\prime}}=\langle n^{\prime}|\rho|n^{\prime}\rangle$. The master equation (II.2) can be simplified into a balance equation Dyk06
$\frac{dp_{n^{\prime}}}{dt}=\kappa\sum_{m^{\prime}}(W_{n^{\prime},m^{\prime}}p_{m^{\prime}}-W_{m^{\prime},n^{\prime}}p_{n^{\prime}}),$ (31)
where the transition rate from level $|m^{\prime}\rangle$ to level $|n^{\prime}\rangle$ ($m^{\prime}\neq n^{\prime}$) is given by
$W_{n^{\prime},m^{\prime}}$ $=$ $M\langle m^{\prime}|a|n^{\prime}\rangle^{*}\langle n^{\prime}|a|m^{\prime}\rangle^{*}+M^{*}\langle m^{\prime}|a|n^{\prime}\rangle\langle n^{\prime}|a|m^{\prime}\rangle$ (32)
$+(1+\bar{N})|\langle n^{\prime}|a|m^{\prime}\rangle|^{2}+\bar{N}|\langle m^{\prime}|a|n^{\prime}\rangle|^{2}.$
One can prove that the transition rate $W_{n^{\prime},m^{\prime}}$ for $n^{\prime}\neq m^{\prime}$ is equal to $|\langle N|a|M\rangle|^{2}$ calculated above in Eq. (III.1.2) .
As can be seen from the above rate equation, even at zero temperature, the oscillator can make transitions to both lower and higher energy levels. In Fig. 4, we compare the stationary distribution obtained using our double perturbation theory with exact numerical results. To the lowest order, all superposition coefficients are zero. To the first order of $\sqrt{\lambda}$ (i.e., $\xi_{kn}=\sqrt{\lambda}\xi^{(1)}_{kn}$), there is no correction to the stationary distribution $p_{n^{\prime}}$. Then, to the second order (i.e., $\xi_{kn}=\sqrt{\lambda}\xi^{(1)}_{kn}+\lambda\xi^{(2)}_{kn}$), we solve the balance equation (31) accordingly. Fig.4 (b) illustrates the relative error $\Delta p_{n}/p_{n}$ in comparison to the exact numerical results, which shows that the discrepancy for low levels is mitigated by high-order perturbative calculations.
In the vicinity of the bottom, we can apply the harmonic approximation ($\xi_{kn}=0$). Under this approximation, the ratio of probabilities over adjacent levels is
$\frac{p_{n^{\prime}+1}}{p_{n^{\prime}}}$ $=$ $\frac{W_{n^{\prime}+1,n^{\prime}}}{W_{n^{\prime},n^{\prime}+1}}=\frac{\bar{N}}{1+\bar{N}}$ (33)
$=$ $\frac{\bar{n}|v|^{2}+(1+\bar{n})|u|^{2}}{1+\bar{n}|v|^{2}+(1+\bar{n})|u|^{2}}.$
We verify the relationship between $\ln(p_{1}/p_{2})$ and the Bose distribution $\bar{n}$ in Fig. 5. The agreement between the analytical results from Eq. (33) and the exact numerical one is excellent.
For high levels, we can define the level-dependent effective temperature via
$N_{eff}(n)\equiv p_{n+1}/(p_{n}-p_{n+1})$
for level $|n\rangle$. Fig. 6 illustrates how $N_{eff}$ varies from the lowest to higher levels. The zero-order term yields a constant effective temperature. When we include the correction to the order of $\lambda$, the correction leads to changes in $N_{eff}$ showing a good agreement with the numerical results for levels near the bottom.
Figure 7: Relationship between effective frequency $\omega_{eff}=\Delta E_{1}$ and damping strength $\kappa$. Results obtained from the emission spectrum method (circles), cf. Eq. (34), are compared with those from our perturbation theory (solid and dashed lines).
Figure 8: Renormalized effective Bose distribution $\tilde{N}$ due to dephasing effects for (a) the LAS and (b) the HAS. We compare the theoretical predictions from Eq. (36) with the exact numerical results and extract the linear relationships.
III.2 Strong Damping and Dephasing
In this section, we explore the dynamics and stationary state of the system under conditions of strong damping and dephasing, which can significantly alter the behavior predicted by the weak damping approximation.
III.2.1 Strong damping
In the regime of strong damping, the harmonic approximation, which leads to the level spacing $\Delta E_{n}$ that is independent of damping, is no longer valid. Instead, the level spacing undergoes slight renormalization for strong damping. Such effect can be observed through the emission spectrum $S(\omega)$, representing the spectral density of photons emitted by the driven resonator, is given by
$S(\omega)=2\operatorname{Re}\int_{0}^{\infty}e^{-i\omega t}\text{Tr}[a^{\dagger}e^{\mathcal{L}t}(a\rho_{\text{st}})]{\rm d}t\,.$ (34)
Our method inherently incorporates damping effects into the calculation of $\Delta E_{n}$. For finite damping, the squeezing parameters $u$ and $v$ are complex numbers, determined by the steady-state conditions given in Eq. (29). Substituting them into Eq. (26), we obtain the effective frequency $\omega_{eff}=\Delta E_{1}$. In Fig. 7, we compare the results obtained through the emission spectrum Eq. (34) with those from our perturbation theory, demonstrating satisfactory consistency. The minor discrepancy arises primarily from higher-level spacings $\Delta E_{n}(n>1)$, which are typically smaller than the first-level spacing $\Delta E_{1}$. For more accurate results, the average of all level spacings should be considered in the calculations.
III.2.2 Dephasing
To incorporate dephasing, we introduce the dephasing term $\eta^{ph}{\cal D}[a^{\dagger}a]{\rho}$ into the master equation (7). For convenience, we define a generalized Lindblad operator ${\cal L}[A;B]\rho\equiv 2A\rho B-BA\rho-\rho AB$. In the spirit of the rotating wave approximation, we obtain the renormalized master equation for the displaced and squeezed density ioperator $\bar{\rho}=S^{\dagger}D^{\dagger}{\rho}DS$ (see detailed derivation in Appendix B)
$\frac{d\bar{\rho}}{dt}$ $=$ $-i[\bar{H},\bar{\rho}]+\frac{\kappa}{2}\{(1+\tilde{N}){\cal D}[a]\bar{\rho}+\tilde{N}{\cal D}[a^{\dagger}]\bar{\rho}\}$ (35)
$+\frac{\kappa}{2}M{\cal L}[a^{\dagger};a^{\dagger}]{\bar{\rho}}+\frac{\kappa}{2}M^{*}{\cal L}[a;a]{\bar{\rho}}$
$+\eta^{ph}(|v|^{2}+|u|^{2})^{2}{\cal D}[a^{\dagger}a]{\bar{\rho}}$
$+\eta^{ph}|uv|^{2}({\cal D}[{a^{\dagger}}^{2}]{\bar{\rho}}+{\cal D}[a^{2}]{\bar{\rho}}),$
where the renormalized Bose distribution, affected by dephasing, is given by:
$\tilde{N}$ $=$ $\bar{N}+\frac{\eta^{ph}}{\kappa}|\alpha^{*}u+\alpha v^{*}|^{2}$ (36)
$=$ $|u|^{2}+\bar{n}(|u|^{2}+|v|^{2})+\frac{\eta^{ph}}{\kappa}|\alpha^{*}u+\alpha v^{*}|^{2}\,.$
We verify our predictions for renormalized Bose distribution by comparing them with exact numerical simulations according to the probabilities over the two lowest levels, specifically $\tilde{N}=p_{2}/(p_{1}-p_{2})$. In Fig. 8, we plot and extract the renormalized Bose distribution as a function of the dephasing, which demonstrates an excellent agreement between the numerical results and the analytical predictions given by Eq. (36) for both LAS and HAS.
IV Conclusions
In this work, we have investigated the quantum dissipative dynamics of a driven Duffing oscillator near the bottoms of its stable states. We elucidated the intricate interplay between the nonlinearity, quantum fluctuations, and the influence of an external driving field. We formulated an effective quantum master equation that encompasses quantum and thermal fluctuations, strong damping, and dephasing within a unified framework. We have developed a refined perturbation theory to analyze the quantum dynamics near both the LAS and the HAS of the Duffing oscillator. While the LAS behavior can be approximated using a harmonic oscillator model, the HAS exhibits more complex behavior due to the significant nonlinear terms. Due to quantum fluctuations, even at zero temperature, higher energy levels near the bottom of the potential well are excited. We calculated the level spacing and effective temperature near the bottom of the HAS and compare them with numerical simulations, demonstrating the accuracy and utility of our proposed approach.
We also investigated the effects of strong damping and dephasing on the system’s dynamics. We showed that the level spacing undergoes slight renormalization for strong damping, which can be observed through the emission spectrum. We derived the renormalized quantum master equation and analyzed the system’s behavior affected by dephasing. Our work provides new insights into the quantum dynamics of driven Duffing oscillators, particularly near their stable states, and offers a theoretical framework that can be applied to related quantum systems under strong damping and dephasing conditions.
ACKNOWLEDGMENTS
This work was supported by Natural Science Foundation of China (Grant No. 12475025).
AUTHOR DECLARATIONS
Conflict of interest
The authors have no conflicts to disclose.
Author contributions
The two authors contributed equally to this work.
DATA AVAILABILITY
The data that support the findings of this study are available within the article.
Appendix A Double Perturbation Theory
To effectively address Hamiltonians containing second-order perturbation terms, exemplified by the HAS in the driven Duffing oscillator, we develop a framework for double perturbation theory. This theory is specifically designed for systems where treating the second-order terms independently is essential for maintaining computational precision and obtaining physically meaningful results. Consider a Hamiltonian of the general form:
${H}=H_{0}+\gamma H_{1}+\gamma^{2}H_{2},$ (37)
where $\gamma$ is a small parameter. In conventional perturbation theory, the terms $\gamma H_{1}+\gamma^{2}H_{2}$ are often treated as a single perturbation term. However, for the HAS of the driven Duffing system, this approach does not yield results with the necessary accuracy. Therefore, we introduce the concept of double perturbation theory, where the terms are handled separately.
The eigenvalues and eigenstates of $H_{0}$ are denoted as $\epsilon^{(0)}_{n}$ and $|\psi^{(0)}_{n}\rangle$ respectively, satisfying $H_{0}|\psi^{(0)}_{n}\rangle=\epsilon^{(0)}_{n}|\psi^{(0)}_{n}\rangle$. The exact eigenvalues and eigenstates of $H$ are denoted as $\epsilon_{n}$ and $|\psi_{n}\rangle$, which can be expanded in powers of $\gamma$ as:
$\epsilon_{n}$ $=$ $\epsilon^{(0)}_{n}+\gamma\epsilon^{(1)}_{n}+\gamma^{2}\epsilon^{(2)}_{n}+\gamma^{3}\epsilon^{(3)}_{n}+\gamma^{4}\epsilon^{(4)}_{n}+o(\gamma^{5}),$
$|\psi_{n}\rangle$ $=$ $|\psi^{(0)}_{n}\rangle+\gamma\sum_{k\neq n}\xi^{(1)}_{kn}|\psi^{(0)}_{k}\rangle+\gamma^{2}\sum_{l\neq n}\xi^{(2)}_{ln}|\psi^{(0)}_{l}\rangle$
$+\gamma^{3}\sum_{m\neq n}\xi^{(3)}_{mn}|\psi^{(0)}_{m}\rangle+\gamma^{4}\sum_{p\neq n}\xi^{(4)}_{pn}|\psi^{(0)}_{p}\rangle+o(\gamma^{5}).$
From the eigenvalue equation $H|\psi_{n}\rangle=\epsilon_{n}|\psi_{n}\rangle$, we can derive the perturbative results order by order:
1) to the order of $\gamma^{0}=1$: $H_{0}|\psi^{(0)}_{n}\rangle=\epsilon^{(0)}_{n}|\psi^{(0)}_{n}\rangle$;
2) to the order of $\gamma$:
$\gamma(\sum_{k\neq n}\epsilon^{(0)}_{k}\xi^{(1)}_{kn}|\psi^{(0)}_{k}\rangle+H_{1}|\psi^{(0)}_{n}\rangle)$ (39)
$=$ $\gamma(\sum_{k\neq n}\epsilon^{(0)}_{n}\xi^{(1)}_{kn}|\psi^{(0)}_{k}\rangle+\epsilon^{(1)}_{n}|\psi^{(0)}_{n}\rangle)\,,$
which gives us the perturbative result to the first order
$\epsilon^{(1)}_{n}$ $=$ $\langle\psi^{(0)}_{n}|H_{1}|\psi^{(0)}_{n}\rangle,$
$\xi^{(1)}_{kn}$ $=$ $\frac{1}{\epsilon^{(0)}_{n}-\epsilon^{(0)}_{k}}\langle\psi^{(0)}_{k}|H_{1}|\psi^{(0)}_{n}\rangle;$ (40)
3) to the order of $\gamma^{2}$:
$\gamma^{2}(\sum_{l\neq n}\epsilon^{(0)}_{l}\xi^{(2)}_{ln}|\psi^{(0)}_{l}\rangle+\sum_{k\neq n}\xi^{(1)}_{kn}H_{1}|\psi^{(0)}_{k}\rangle+H_{2}|\psi^{(0)}_{n}\rangle)$
$=\gamma^{2}(\sum_{l\neq n}\epsilon^{(0)}_{n}\xi^{(2)}_{ln}|\psi^{(0)}_{l}\rangle+\sum_{k\neq n}\xi^{(1)}_{kn}\epsilon^{(1)}_{n}|\psi^{(0)}_{k}\rangle+\epsilon^{(2)}_{n}|\psi^{(0)}_{n}\rangle),$
which gives us the second order perturbative result:
$\epsilon^{(2)}_{n}$ $=$ $\sum_{k\neq n}\langle\psi^{(0)}_{n}|H_{1}|\psi^{(0)}_{k}\rangle\xi^{(1)}_{kn}+\langle\psi^{(0)}_{n}|H_{2}|\psi^{(0)}_{n}\rangle,$
$\xi^{(2)}_{ln}$ $=$ $\frac{1}{\epsilon^{(0)}_{n}-\epsilon^{(0)}_{l}}(\sum_{k\neq n}\langle\psi^{(0)}_{l}|H_{1}|\psi^{(0)}_{k}\rangle\xi^{(1)}_{kn}+\langle\psi^{(0)}_{l}|H_{2}|\psi^{(0)}_{n}\rangle$ (42)
$-\epsilon^{(1)}_{n}\xi^{(1)}_{ln});$
4) to the order of $\gamma^{3}$:
$\gamma^{3}(\sum_{m\neq n}\epsilon^{(0)}_{m}\xi^{(3)}_{mn}|\psi^{(0)}_{m}\rangle+\sum_{l\neq n}\xi^{(2)}_{ln}H_{1}|\psi^{(0)}_{l}\rangle$ (43)
$+\sum_{k\neq n}\xi^{(1)}_{kn}H_{2}|\psi^{(0)}_{k}\rangle)$
$=$ $\gamma^{3}(\sum_{m\neq n}\epsilon^{(0)}_{n}\xi^{(3)}_{mn}|\psi^{(0)}_{m}\rangle+\sum_{l\neq n}\epsilon^{(1)}_{n}\xi^{(2)}_{ln}|\psi^{(0)}_{l}\rangle$
$+\sum_{k\neq n}\xi^{(1)}_{kn}\epsilon^{(2)}_{n}|\psi^{(0)}_{k}\rangle+\epsilon^{(3)}_{n}|\psi^{(0)}_{n}\rangle),$
which gives us the perturbative result to the third order:
$\epsilon^{(3)}_{n}$ $=$ $\sum_{l\neq n}(\langle\psi^{(0)}_{n}|H_{1}|\psi^{(0)}_{l}\rangle\xi^{(2)}_{ln}+\langle\psi^{(0)}_{n}|H_{2}|\psi^{(0)}_{l}\rangle\xi^{(1)}_{ln})\,,$
$\xi^{(3)}_{mn}$ $=$ $\frac{1}{\epsilon^{(0)}_{n}-\epsilon^{(0)}_{m}}[\sum_{l\neq n}(\langle\psi^{(0)}_{m}|H_{1}|\psi^{(0)}_{l}\rangle\xi^{(2)}_{ln}$ (44)
$+\langle\psi^{(0)}_{m}|H_{2}|\psi^{(0)}_{l}\rangle\xi^{(1)}_{ln})-\epsilon^{(1)}_{n}\xi^{(2)}_{mn}-\epsilon^{(2)}_{n}\xi^{(1)}_{mn}];$
5) to the order of $\gamma^{4}$:
$\gamma^{4}(\sum_{p\neq n}\epsilon^{(0)}_{p}\xi^{(4)}_{pn}|\psi^{(0)}_{p}\rangle+\sum_{m\neq n}\xi^{(3)}_{mn}H_{1}|\psi^{(0)}_{m}\rangle+\sum_{l\neq n}\xi^{(2)}_{ln}H_{2}|\psi^{(0)}_{l}\rangle)$
$=\gamma^{4}[\sum_{m\neq n}(\epsilon^{(0)}_{n}\xi^{(4)}_{mn}+\epsilon^{(1)}_{n}\xi^{(3)}_{mn}+\xi^{(2)}_{mn}\epsilon^{(2)}_{n}+\xi^{(1)}_{mn}\epsilon^{(3)}_{n})|\psi^{(0)}_{m}\rangle$
$+\epsilon^{(4)}_{n}|\psi^{(0)}_{n}\rangle],$ (45)
which gives us the fourth order perturbative result:
$\epsilon^{(4)}_{n}$ $=$ $\sum_{l\neq n}(\langle\psi^{(0)}_{n}|H_{1}|\psi^{(0)}_{l}\rangle\xi^{(3)}_{ln}+\langle\psi^{(0)}_{n}|H_{2}|\psi^{(0)}_{l}\rangle\xi^{(2)}_{ln}),$
$\xi^{(4)}_{pn}$ $=$ $\frac{1}{\epsilon^{(0)}_{n}-\epsilon^{(0)}_{p}}[\sum_{l\neq n}(\langle\psi^{(0)}_{p}|H_{1}|\psi^{(0)}_{l}\rangle\xi^{(3)}_{ln}+\langle\psi^{(0)}_{p}|H_{2}|\psi^{(0)}_{l}\rangle\xi^{(2)}_{ln})$ (46)
$-\epsilon^{(1)}_{n}\xi^{(3)}_{pn}-\epsilon^{(2)}_{n}\xi^{(2)}_{pn}-\epsilon^{(3)}_{n}\xi^{(1)}_{pn}].$
Similar expressions can be derived for higher-order corrections.
For the high-amplitude state of the driven Duffing oscillator, the perturbative Hamiltonian is given by Eq. (26),
$\frac{\bar{H}^{h}}{\delta\omega}$ $=$ $h^{h}_{0}+\gamma h^{h}_{1}+\gamma h^{h}_{2},$
where the order parameter $\gamma$ is $\sqrt{\lambda}$. The zeroth-order Hamiltonian
$h^{h}_{0}$ $=$ $[(1-4\lambda|\alpha_{h}|^{2})(|v|^{2}+|u|^{2})$ (48)
$-2\lambda({\alpha^{*}_{h}}^{2}uv+\alpha_{h}^{2}u^{*}v^{*})]a^{\dagger}a,$
corresponds to a harmonic oscillator, and the eigenstates $|\psi^{(0)}_{k}\rangle$ are simply the harmonic oscillator states $|k\rangle$ with eigenvalues:
$\epsilon^{(0)}_{k}$ $=$ $k[(1-4\lambda|\alpha_{h}|^{2})(|v|^{2}+|u|^{2})$ (49)
$-2\lambda({\alpha^{*}_{h}}^{2}uv+\alpha_{h}^{2}u^{*}v^{*})].$
By calculating the matrix elements $\langle l|h^{h}_{1}|k\rangle$ and $\langle l|h^{h}_{2}|k\rangle$,
$\langle l|h^{h}_{1}|k\rangle$ $=$ $-2\sqrt{\lambda k}[\alpha_{h}|v|^{2}u^{*}(2k-1)+\alpha_{h}|u|^{2}u^{*}(k+1)+\alpha_{h}^{*}v|u|^{2}(2k+1)+\alpha_{h}^{*}v|v|^{2}(k-1)+(\alpha_{h}^{*}v+\alpha_{h}u^{*})/2]\delta_{l,k-1}$
$-2\sqrt{\lambda(k+1)}[\alpha_{h}|u|^{2}v^{*}(2k+3)+\alpha_{h}|v|^{2}v^{*}k+\alpha_{h}^{*}u|v|^{2}(2k+1)+\alpha_{h}^{*}u|u|^{2}(k+2)+(\alpha_{h}^{*}u+\alpha_{h}v^{*})/2]\delta_{l,k+1}$
$-2u^{*}v\sqrt{\lambda k(k-1)(k-2)}(\alpha_{h}^{*}v+\alpha_{h}u^{*})\delta_{l,k-3}-2v^{*}u\sqrt{\lambda(k+1)(k+2)(k+3)}(\alpha_{h}^{*}u+\alpha_{h}v^{*})\delta_{l,k+3},$
$\langle l|h^{h}_{2}|k\rangle$ $=$ $-[2(2|uv|^{2}+|u|^{4})k+(|u|^{4}+|v|^{4}+4|uv|^{2})k^{2}]\delta_{l,k}$ (50)
$-2u^{*}v\sqrt{k(k-1)}[k|u|^{2}+(k-1)|v|^{2}]\delta_{l,k-2}-2v^{*}u\sqrt{(k+1)(k+2)}[(k+2)|u|^{2}+(k+1)|v|^{2}]\delta_{l,k+2}$
$-(u^{*}v)^{2}\sqrt{k(k-1)(k-2)(k-3)}\delta_{l,k-4}-(v^{*}u)^{2}\sqrt{(k+1)(k+2)(k+3)(k+4)}\delta_{l,k+4},$
we can apply the double perturbation theory to obtain the desired perturbative results, such as level spacing and effective temperature, for the high-amplitude state as detailed in the main text.
Appendix B Dephasing
This section presents the derivation of Eq. (35) in the main text. To incorporate dephasing, we introduce the dephasing term $\eta^{ph}{\cal D}[a^{\dagger}a]{\rho}$ into the master equation. For convenience, we define a generalized Lindblad operator ${\cal L}[A;B]\rho\equiv 2A\rho B-BA\rho-\rho AB$, which satisfies
${\cal D}[A]{\rho}={\cal L}[A;A^{\dagger}]\rho,$
${\cal D}[A+B]{\rho}={\cal D}[A]{\rho}+{\cal D}[B]{\rho}+{\cal L}[A;B^{\dagger}]\rho+{\cal L}[B;A^{\dagger}]\rho,$
${\cal L}[A;B+C]\rho={\cal L}[A;B]\rho+{\cal L}[A;C]\rho,$
${\cal L}[A+B;C]\rho={\cal L}[A;C]\rho+{\cal L}[B;C]\rho,$
$({\cal L}[A;B]\rho)^{\dagger}={\cal L}[B^{\dagger};A^{\dagger}]\rho.$ (51)
Utilizing these properties, we derive the transformed dephasing term under the action of the displacement and squeezing operators:
$S^{\dagger}D^{\dagger}({\cal D}[a^{\dagger}a]{\rho})DS$
$=(|v|^{2}+|u|^{2})^{2}{\cal D}[a^{\dagger}a]{\bar{\rho}}\ +|\alpha^{*}u+\alpha v^{*}|^{2}({\cal D}[a^{\dagger}]{\bar{\rho}}$
$+{\cal D}[a]{\bar{\rho}})+|uv|^{2}({\cal D}[{a^{\dagger}}^{2}]{\bar{\rho}}+{\cal D}[a^{2}]{\bar{\rho}})+{\cal L}\bar{\rho}+({\cal L}\bar{\rho})^{\dagger}.$ (52)
The term ${\cal L}\bar{\rho}$ includes various higher-order terms:
${\cal L}\bar{\rho}$ $=$ $|uv|^{2}{\cal L}[{a^{\dagger}}^{2};{a^{\dagger}}^{2}]{\bar{\rho}}+|\alpha^{*}u+\alpha v^{*}|^{2}{\cal L}[{a^{\dagger}};{a^{\dagger}}]{\bar{\rho}}+v^{*}u(\alpha^{*}u+\alpha v^{*}){\cal L}[{a^{\dagger}}^{2};{a^{\dagger}}]{\bar{\rho}}+vu^{*}(\alpha u^{*}+\alpha^{*}v){\cal L}[a;a^{2}]{\bar{\rho}}$
$+v^{*}u(\alpha u^{*}+\alpha^{*}v){\cal L}[a;{a^{\dagger}}^{2}]{\bar{\rho}}+vu^{*}(\alpha^{*}u+\alpha v^{*}){\cal L}[a^{2};a^{\dagger}]{\bar{\rho}}+(|v|^{2}+|u|^{2})\{v^{*}u{\cal L}[a^{\dagger}a;{a^{\dagger}}^{2}]{\bar{\rho}}+vu^{*}{\cal L}[a^{\dagger}a;a^{2}]{\bar{\rho}}$
$+(\alpha^{*}u+\alpha v^{*}){\cal L}[a^{\dagger};a^{\dagger}a]{\bar{\rho}}+(\alpha u^{*}+\alpha^{*}v){\cal L}[a;a^{\dagger}a]{\bar{\rho}}\}.$
In the spirit of the rotating wave approximation, we neglect the terms ${\cal L}\bar{\rho}$ and $({\cal L}\bar{\rho})^{\dagger}$ in Eq. (B). Adding the dephasing term to the renormalized master equation, we obtain:
$\frac{d\bar{\rho}}{dt}$ $=$ $-i[\bar{H},\bar{\rho}]+\frac{\kappa}{2}\{(1+\tilde{N}){\cal D}[a]\bar{\rho}+\tilde{N}{\cal D}[a^{\dagger}]\bar{\rho}\}+\frac{\kappa}{2}M{\cal L}[a^{\dagger};a^{\dagger}]{\bar{\rho}}+\frac{\kappa}{2}M^{*}{\cal L}[a;a]{\bar{\rho}}$ (54)
$+\eta^{ph}(|v|^{2}+|u|^{2})^{2}{\cal D}[a^{\dagger}a]{\bar{\rho}}+\eta^{ph}|uv|^{2}({\cal D}[{a^{\dagger}}^{2}]{\bar{\rho}}+{\cal D}[a^{2}]{\bar{\rho}}),$
where the renormalized Bose distribution, affected by dephasing, is given by:
$\tilde{N}$ $=$ $\bar{N}+\frac{\eta^{ph}}{\kappa}|\alpha^{*}u+\alpha v^{*}|^{2}$ (55)
$=$ $|u|^{2}+\bar{n}(|u|^{2}+|v|^{2})+\frac{\eta^{ph}}{\kappa}|\alpha^{*}u+\alpha v^{*}|^{2}\,.$

### v2

@@
 We investigate the quantum dissipative dynamics near the stable states (attractors) of a driven Duffing oscillator.
-A refined perturbation theory that can treat two perturbative parameters with different orders is developed to calculate the quantum properties of Duffing oscillator near the attractors.
-We obtain the perturbative analytical results for the renormalized level spacings and the quantum-fluctuation-induced distribution of higher energy levels near the attractor, which are then verified by numerical simulations.
-Furthermore, we demonstrate that strong damping induces a slight renormalization of level spacing, and the renormalized Bose distribution affected by dephasing.
-Our work provides new insights into the quantum dynamics of driven Duffing oscillator and offers a theoretical framework that can be applied to related quantum system near their stable states.
+A refined perturbation theory that can treat two perturbative parameters with different orders is developed to calculate the quantum properties of the Duffing oscillator near the attractors.
+We obtain the perturbative analytical results, that go beyond the standard linearization approach, for the renormalized level spacings, the orbital displacements, and the effective temperature near the classical attractor.
+Furthermore, we demonstrate that strong damping induces additional slight renormalization of level spacings, and the Bose distribution together with dephasing.
+Our work provides new insights into the quantum dynamics of the driven Duffing oscillator and offers a theoretical framework that can be applied to related quantum systems near their stable states.
 I Introduction
-The driven Duffing oscillator Duf18 , a paradigmatic model for various nonlinear mechanical systems and nonlinear optical phenomena, has been fascinating physicists for a long time by its rich dynamical behaviors such as bistability, bifurcation, and chaotic trajectories.
+The driven Duffing oscillator Duf18 , a paradigmatic model for various nonlinear mechanical systems and nonlinear optical phenomena, has fascinated physicists for a long time with its rich dynamical behaviors such as bistability, bifurcation, and chaotic trajectories.
 In recent years, the Duffing oscillator has received renewed attention as the quantum regime of nanomechanical oscillators becomes experimentally accessible.
-The interplay between quantum effects, nonequilibrium dynamics, and nonlinear effects makes the driven Duffing oscillator a model with broad applicability across multiple domains, e.g., mechanical metrology Dyk12m ; Cle02 ; Lif08 ; Poo12 , cavity and circuit quantum electrodynamics Jea24 ; Mav16 ; Boi10 , nano- and opto-mechanics Buk06 ; Alm07 ; Del07 ; Kip11 ; Dyk22 ; Pet14 ; Bol23 ; Hel24 ; Dyk22R , and cold atoms Art03 ; Got19 ; Par12 .
-Notable examples include bifurcation-based quantum measurement devices, where the low- and high-amplitude states of bistability are entangled with the ground and excited states of qubits respectively, enabling the analysis of the qubit’s state through the detection of classical signals Ald05 ; Sid04 ; Sid06 ; Man07 ; Sid09 .
-The dissipative dynamics of driven Duffing oscillators have been extensively studied.
+The interplay between quantum effects, nonequilibrium dynamics, and nonlinear effects makes the driven Duffing oscillator a model with broad applicability across multiple domains, e.g., mechanical metrology Dyk12m ; Cle02 ; Lif08 ; Poo12 , chaotic dynamics Kro23 ; Gho23 , cavity and circuit quantum electrodynamics Zhang13n ; Jea24 ; Mav16 ; Boi10 , nano- and opto-mechanics Buk06 ; Alm07 ; Del07 ; Kip11 ; Dyk22 ; Ken24 ; Pet14 ; Bol23 ; Hel24 ; Dyk22R ; Mil08 , and cold atoms Art03 ; Got19 ; Par12 .
+Notable examples include bifurcation-based quantum measurement devices, where the low- and high-amplitude states of bistability are entangled with the ground and excited states of qubits, respectively, enabling the analysis of the qubit states through the detection of classical signals Ald05 ; Sid04 ; Sid06 ; Man07 ; Sid09 .
+In practice, physical systems inevitably interact with their environment, leading to the decoherence of quantum states and dissipation of energy.
+The dissipative dynamics of driven Duffing oscillators have been extensively studied Chen23New ; Mar21New .
 In the underdamped regime near a bifurcation point, a scaling law for the noise-induced escape from metastable states was established Dyk05 .
+Close to the bottom of the well in a parametrically driven Duffing oscillator, it was identified that the energy dependence of the level spacings captured by the perturbative approach beyond linearization gives rise to a fine structure in the power spectrum Dyk11n .
 It was also revealed that the quantum activation process has distinct temperature dependency compared to that for the quantum tunneling process Dyk07 .
@@
 We develop an effective quantum master equation that can address quantum fluctuations, thermal effects, damping, and dephasing in a unified framework.
-We quantify the occupation of high levels near the bottom of the potential well by the the quantum squeezing effects.
+Our approach is essential to quantify the occupation of high levels near the bottom of the potential well by the quantum squeezing effects.
 We also demonstrate the effects of strong damping and dephasing on the system’s dynamics, including level spacing renormalization and dephasing-modified Bose distributions.
+While prior studies within linearized frameworks have successfully captured phenomena such as effective temperatures Dyk11n ; Pea10n ; Lem15n , the refined perturbation theory presented in our work goes beyond the standard linearization approach allowing us to investigate the energy dependence of the level spacing, the orbital displacement, and the effective temperature.
+To address this, we develop a refined perturbation theory capable of treating two perturbative parameters with distinct orders, enabling a unified analysis of nonlinear and dissipative quantum effects near the attractors.
 By comparing our theoretical predictions with exact numerical simulations, we demonstrate the accuracy and utility of our proposed framework.
 II General Theory
-Figure 1: (a) Quasienergy landscape in phase space of a driven Duffing oscillator.
-The extrema correspond to the high amplitude stable state (M) and the low amplitude stable state (m), while the saddle point (s) denotes an unstable state.
-(b) Cross-section of the quasienergy potential at $P=0$.
-Quantum energy levels close to the maximum (M) and minimum (m) of the potential are depicted by blue and red lines, respectively.
-The unstable saddle point (s) is also indicated.
 II.1 Model Hamiltonian
-An extensive class of macroscopic physical systems, such as Josephson junctions and nanomechanical oscillators can be modeled by the Duffing oscillator in the presence of a periodic driving force with the system Hamiltonian described by Ser07
+An extensive class of macroscopic physical systems, such as Josephson junctions and nanomechanical oscillators can be modeled by the Duffing oscillator in the presence of a periodic driving force, with the system Hamiltonian described by Ser07
 ${H}_{S}(t)=\frac{{p}^{2}}{2m}+\frac{1}{2}m\Omega^{2}{x}^{2}-\gamma{x}^{4}+F(t){x}.$ (1)
-Here, parameter $m$ ($\Omega$) describes the mass (frequency) of the oscillator, $\gamma$ gives the nonlinearity of Duffing oscillator, and $F(t)=F_{0}(e^{i\nu t}+e^{-i\nu t})$ describes the periodically driving force with frequency $\nu$.
-By switching to the rotating frame using the transformation $U(t)=\exp(-i\nu a^{\dagger}at)$, where $a^{\dagger}$ and $a$ are the oscillator raising and lowering operators, and applying the rotating wave approximation (RWA), we obtain a time-independent Hamiltonian
-$H=(\delta\omega+\chi)a^{\dagger}a+\chi(a^{\dagger}a)^{2}+\epsilon(a^{\dagger}+a)\,,$ (2)
-where $\delta\omega=\hbar(\Omega-\nu)$ is the frequency detuning, $\chi=-3\gamma\hbar^{2}/2(m\Omega)^{2}$ is the scaled dimensionless nonlinearity, and $\epsilon=F_{0}\sqrt{{\hbar}/{2m\Omega}}$ is the scaled driving strength.
+Here, parameter $m$ ($\Omega$) describes the mass (frequency) of the oscillator, $\gamma$ gives the nonlinearity of Duffing oscillator, and $F(t)=F_{0}(e^{i\nu t}+e^{-i\nu t})$ describes the periodical driving force with frequency $\nu$.
+By switching to the rotating frame using the transformation $U(t)=\exp(-i\nu a^{\dagger}at)$ with $a^{\dagger}$ ($a$) the raising (lowering) operator, and applying the rotating wave approximation (RWA), we obtain a time-independent Hamiltonian
+$H=(\delta\omega+\chi)a^{\dagger}a+\chi(a^{\dagger}a)^{2}+\epsilon(a^{\dagger}+a).$ (2)
+Here, parameter $\delta\omega=\hbar(\Omega-\nu)$ is the frequency detuning, $\chi=-3\gamma\hbar^{2}/2(m\Omega)^{2}$ is the scaled dimensionless nonlinearity, and $\epsilon=F_{0}\sqrt{{\hbar}/{2m\Omega}}$ is the scaled driving strength.
 We then introduce the position operator $Q$ and momentum operator $P$ in the rotating frame via
@@
 II.2 Renormalized master equation
+Figure 1: (a) Quasienergy landscape in phase space of a driven Duffing oscillator.
+The extrema correspond to the high-amplitude stable state (M) and the low-amplitude stable state (m), while the saddle point (s) marks an unstable state.
+(b) Cross-section of the quasienergy potential at $P=0$.
+Quantum energy levels close to the maximum (M) and minimum (m) of the potential are depicted by blue and red lines, respectively.
+The unstable saddle point (s) is also indicated.
 The characteristic behavior of the driven Duffing system is the bistability manifesting as two stable states: the low-amplitude state (LAS) and the high-amplitude state (HAS).
@@
 The unstable state, known as the saddle point, is located on the separatrix, serving as the boundary dividing the basins of the attractors.
-When damping is present, the system evolves towards the nearby attractor if its initial state lies within the basin of the attractor.
+When the damping is present, the system evolves towards the nearby attractor if its initial state lies within the basin of the attractor.
 However, due to thermal noise, the system does not remain exactly on the attractor, but forms a probability distribution within its basin of attraction.
-To describe the dissipative dynamics of our system, we employ a Lindblad form master equation
+To describe the dissipative dynamics of our system, we employ the Lindblad form master equation
 $\frac{d\rho}{dt}=-i[H,{\rho}]+\frac{\kappa}{2}\{(1+\bar{n}){\cal D}[a]{\rho}+\bar{n}{\cal D}[a^{\dagger}]{\rho}\},$ (7)
 where ${\cal D}[\bullet]$ is the Lindblad operator defined as ${\cal D}[A]{\rho}\equiv 2A{\rho}A^{\dagger}-A^{\dagger}A{\rho}-{\rho}A^{\dagger}A$, $\bar{n}$ is the Bose-Einstein distribution, and $\kappa$ is the damping strength.
-To study the quantum dynamics near the bottom of each stable state, we first transform the system to the center of the attractor using the displacement operator
+To study the quantum dynamics near the bottom of each stable state,vwe first transform the system to the center of the attractor using the displacement operator
 $D[\alpha]=e^{\alpha a^{\dagger}-\alpha^{*}a}$ (8)
@@
 $+\left[{\tilde{\alpha}}^{*}a-{\tilde{\alpha}}a^{\dagger},\tilde{\rho}\right]\,,$
-with ${\tilde{\alpha}}=[\frac{1}{2}\kappa+i(\delta\omega+\chi+2\chi\alpha|^{2})]\alpha+i\varepsilon.$ By choosing $\alpha$ such that $\tilde{\alpha}=0$, the master equation is simplified into
+with ${\tilde{\alpha}}=[\frac{1}{2}\kappa+i(\delta\omega+\chi+2\chi|\alpha|^{2})]\alpha+i\varepsilon.$ By choosing $\alpha$ such that $\tilde{\alpha}=0$, the master equation is simplified into
 $\frac{d\tilde{\rho}}{dt}$ $=$ $-[\tilde{H},{\tilde{\rho}}]+\frac{\kappa}{2}\{(1+\bar{n}){\cal D}[a]{\tilde{\rho}}+\bar{n}{\cal D}[a^{\dagger}]{\tilde{\rho}}\}$ (10)
@@
 which has the transformation property $S^{\dagger}aS=va+ua^{\dagger}$ with $v=\cosh(|\xi|)$ and $u=-\frac{|\xi|}{\xi}\sinh(|\xi|)$.
-Bu defining the squeezed density operator $\bar{\rho}=S^{\dagger}\tilde{\rho}S$, we transform the master equation Eq.
+By defining the squeezed density operator $\bar{\rho}=S^{\dagger}\tilde{\rho}S$, we transform the master equation Eq.
 (10) into the following form
@@
 II.3 Orders of perturbative parameters
-To perform perturbation calculations for the renormalized Hamiltonian (II.2), we can choose dimensionless Planck constant $\lambda=-\chi/(4\Delta)$ as the natural choice for the perturbation parameter.
+To perform perturbation calculations for the renormalized Hamiltonian (II.2), we can choose the dimensionless Planck constant $\lambda=-\chi/(4\Delta)$ as the natural choice for the perturbation parameter.
 However, since the displacement parameter $\alpha$ is also a function of $\lambda$, it is subtle to properly organize the perturbative terms according to their respective orders.
@@
 The stable state of the driven Duffing oscillator can be approximated as a coherent state $|\alpha\rangle$.
-By applying the variational principle in quantum mechanics, i.e., $\partial_{\alpha}\langle\alpha|H|\alpha\rangle=0$, we obtain two solutions of the steady coherent number: a smaller one $|\alpha_{l}|^{2}\approx\beta/(2\lambda)$ for the LAS and a larger one $|\alpha_{h}|^{2}\approx 1/(2\lambda)$ for the HAS.
+By applying the variational principle in quantum mechanics, i.e., $\partial_{\alpha}\langle\alpha|H|\alpha\rangle=0$, we obtain two solutions for the steady coherent number: a smaller one $|\alpha_{l}|^{2}\approx\beta/(2\lambda)$ for the LAS and a larger one $|\alpha_{h}|^{2}\approx 1/(2\lambda)$ for the HAS.
 For the LAS, we sort the terms in the renormalized Hamiltonian (II.2) as follows
 $\frac{\bar{H}^{l}}{\delta\omega}=h^{l}_{0}+\lambda h^{l}_{\lambda}+\beta h^{l}_{\beta}+\sqrt{\lambda\beta}h^{l}_{\lambda\beta},$ (24)
-with each term given by
+where the four terms on the right-hand side are given by
 $h^{l}_{0}$ $=$ $(|v|^{2}+|u|^{2})a^{\dagger}a,$
@@
 We introduce $\gamma=\sqrt{\lambda}$ as the perturbation parameter and rewrite the Hamiltonian as
-$\frac{\bar{H}^{h}}{\delta\omega}$ $=$ $h^{h}_{0}+\gamma h^{h}_{1}+\gamma h^{h}_{2}$ (26)
-with each term given by
+$\frac{\bar{H}^{h}}{\delta\omega}$ $=$ $h^{h}_{0}+\gamma h^{h}_{1}+\gamma h^{h}_{2},$ (26)
+where the three terms on the right-hand side are
 $h^{h}_{0}$ $=$ $[(1-4\lambda|\alpha_{h}|^{2})(|v|^{2}+|u|^{2})$
@@
 However, for the HAS, the nonlinear term $\chi(a^{\dagger}a)^{2}\approx\chi|\alpha_{h}|^{2}$ becomes prominent, and the oscillator behaves as a highly squeezed coherent state.
-In the following sections, we will apply our perturbative method to calculate the crucial quantities related to the HAS of the driven Duffing oscillator, namely the level spacing, averaged position of energy level, and the effective temperature in the vicinity of the HAS attractor.
+In the following sections, we will apply our perturbative method to calculate the crucial quantities related to the HAS of the driven Duffing oscillator, namely, the level spacings, the orbital displacement, and the effective temperature in the vicinity of the HAS attractor.
 III Results
 III.1 Quantum dynamics of HAS
-The quantum dynamics of the driven Duffing oscillator near the HAS attractor exhibit a rich interplay between nonlinearity, quantum fluctuations, and thermal noise.
+The quantum dynamics of the driven Duffing oscillator near the HAS attractor exhibit a rich interplay among nonlinearity, quantum fluctuations, and thermal noise.
 In this section, we discuss the quantum properties of the HAS using the renormalized master equation combined with a refined perturbation theory.
@@
 The exact results (lowest curve) for level spacing $\Delta E_{n}=|E_{n+1}-E_{n}|$ are compared with the zeroth-order (constant spacing), the second-order (second highest line, $\gamma^{2}=\lambda$) and the fourth-order (third highest line, $\gamma^{4}=\lambda^{2}$) corrections.
-Excellent agreement is observed for low level numbers near the potential well bottom, while higher-order corrections are required for large $n$.
+Excellent agreement is observed for low-level numbers near the potential well bottom, while higher-order corrections are required for large $n$.
 Parameters: $\lambda=0.016$ and $\beta=4/75$.
-The nonlinearity term $\chi(a^{\dagger}a)^{2}$ in the Hamiltonian, having an opposite sign to $\delta\omega$, results in a decrease in the level spacing $\Delta E_{n}=|E_{n+1}-E_{n}|$ as we approach the saddle point, as illustrated in Fig.
+The nonlinear term $\chi(a^{\dagger}a)^{2}$ in the Hamiltonian, having the opposite sign to $\delta\omega$, results in a decrease in the level spacing $\Delta E_{n}=|E_{n+1}-E_{n}|$ as we approach the saddle point, as illustrated in Fig.
 1(b).
@@
 (26) as one perturbation term.
-However, it becomes a challenge to control the accuracy of the level spacings by the perturbative parameter.
+However, it becomes a challenge to control the accuracy of the level spacings using the perturbative parameter.
 We find it necessary to distinguish between these two perturbative terms in the perturbation calculations to accurately determine the level spacing.
-To address this, we have developed a double perturbation theory framework that is particularly suited for the HAS Hamiltonian containing second-order small terms, see the details in Appendix A,
+To address this, we have developed a double perturbation theory framework that is particularly suited for the HAS Hamiltonian containing second-order small terms, see the details in Appendix A.
 In Fig.
 2, we compare our perturbation calculations with the exact numerical results obtained by diagonalizing the original Hamiltonian Eq.
-(2), which shows an excellent agreement for energy levels near the bottom of the potential well.
-Under the zeroth-order perturbation approximation, the energy level spacing remains constant across all levels, similar to that of harmonic oscillators.
-The second-order and fourth-order corrections correspond to accuracies up to $\gamma^{2}=\lambda$ and $\gamma^{4}=\lambda^{2}$, respectively.
+(2).
+These results show an excellent agreement for energy levels near the bottom of the potential well.
+Under the zeroth-order perturbation approximation, the energy level spacing remains constant across all levels, similar to that of the harmonic oscillator.
+The second-order and fourth-order corrections provide accuracies up to $\gamma^{2}=\lambda$ and $\gamma^{4}=\lambda^{2}$, respectively.
 Higher-order perturbation calculations become necessary for levels farther from the bottom.
-Figure 3: Average position of energy levels.
-Average position $\langle N|a|N\rangle$ of the energy level $N$, illustrating the shifts due to quantum fluctuations.
+Figure 3: Orbitial displacement.
+The average position $\langle N|a|N\rangle$ of the energy level $N$, illustrates the shift due to the quantum fluctuation.
 Under the harmonic approximation, $\langle N|a|N\rangle$ is a constant for every level (lowest line).
-Considering higher-order corrections, we observe changes in $\langle N|a|N\rangle$ with levels.
-Comparing with the second order perturbation result (highest line), the third order result (second highest line) already align well with the exact one (the bold line).
-III.1.2 Averaged level position
-We denote the eigenstate of the renormalized Hamiltonian $\bar{H}$ as $|n^{\prime}\rangle$ that is generally a superposition of harmonic oscillator eigenstates $|n^{\prime}\rangle=|n\rangle+\sum_{k\neq n}\xi_{kn}|k\rangle$, where the superposition coefficients $\xi_{kn}$ are provided in the Appendix A.
+Considering higher-order corrections, we observe changes in $\langle N|a|N\rangle$ across levels.
+Compared with the second-order perturbation result (highest line), the third-order result (second highest line) already aligns well with the exact one (bold line).
+III.1.2 Orbital displacement
+We denote the eigenstate of the renormalized Hamiltonian $\bar{H}$ as $|n^{\prime}\rangle$, which is generally a superposition of harmonic oscillator eigenstates $|n^{\prime}\rangle=|n\rangle+\sum_{k\neq n}\xi_{kn}|k\rangle$, where the superposition coefficients $\xi_{kn}$ are provided in the Appendix A.
 The eigenstate $|N\rangle$ of the original Hamiltonian (2) is related to that of the renormalized Hamiltonian (II.2) via the relationship $|N\rangle=DS|n^{\prime}\rangle$.
@@
 $=$ $v\langle n^{\prime}|a|m^{\prime}\rangle+u\langle m^{\prime}|a|n^{\prime}\rangle^{*}+\alpha\langle n^{\prime}|m^{\prime}\rangle\,.$
-The matrix element $\langle N|a|N\rangle$ provides insight into the average position of each level in the phase space.
-Under the harmonic approximation ($\xi_{kn}=0$), $\langle N|a|N\rangle$ is a constant $\alpha$ for all levels.
-However, considering higher-order corrections, the average position $\langle N|a|N\rangle$ changes with energy level, as depicted in Fig.
+The matrix element $\langle N|a|N\rangle$ provides insight into the orbital displacement in the phase space.
+Under the harmonic approximation ($\xi_{kn}=0$), $\langle N|a|N\rangle$ remains a constant $\alpha$ for all levels.
+However, considering higher-order corrections, the orbital displacement $\langle N|a|N\rangle$ changes with energy level, as depicted in Fig.
 3.
-The perturbation calculations agree well with numerical results.
+The perturbation results agree well with numerical calculations.
 Figure 4: Comparison of stationary probability distributions.
-(a) Comparison of the stationary probability distribution obtained through perturbation theory to zeroth order (red line), second order (blue lines) and exact numerical simulations (black line).
+(a) Comparison of the stationary probability distribution obtained through perturbation theory to zeroth-order (red line), second-order (blue lines) and exact numerical simulations (black line).
 (b) Relative error $\Delta p_{n}/p_{n}$ in the stationary distribution obtained from perturbation theory compared to numerical results.
@@
 The zero-order effective temperature term gives a constant effective temperature.
-The second order results in changes in $N_{eff}(n)$, which is quite accurate for levels near the bottom.
+The second-order term results in changes in $N_{eff}(n)$, which are quite accurate for levels near the bottom.
 III.1.3 Effective temperature
@@
 It is important to note that the annihilation operator $a$ is for the Fock state $|N\rangle$, which decreases the Fock state from a higher level to the next lower level $a|N\rangle=\sqrt{N}|N-1\rangle$.
-However, in our case, the eigenstate of quasienergy $|n\rangle$ is the superposition of the Fock state $|N\rangle$.
+However, in our case, the eigenstate of quasienergy $|n\rangle$ is the superposition of Fock states $|N\rangle$.
 As a result, the annihilation operator $a$ can either decrease or increase the state $|n\rangle$ even at zero temperature.
@@
 Thus, we can only keep the diagonal elements.
-Here, we assume that the stationary density matrix is diagonal and denote the diagonal terms $p_{n^{\prime}}=\langle n^{\prime}|\rho|n^{\prime}\rangle$.
+Here, we assume that the stationary density matrix is diagonal and denote the diagonal terms as $p_{n^{\prime}}=\langle n^{\prime}|\rho|n^{\prime}\rangle$.
 The master equation (II.2) can be simplified into a balance equation Dyk06
@@
 Then, to the second order (i.e., $\xi_{kn}=\sqrt{\lambda}\xi^{(1)}_{kn}+\lambda\xi^{(2)}_{kn}$), we solve the balance equation (31) accordingly.
-Fig.4 (b) illustrates the relative error $\Delta p_{n}/p_{n}$ in comparison to the exact numerical results, which shows that the discrepancy for low levels is mitigated by high-order perturbative calculations.
+Fig.
+4 (b) illustrates the relative error $\Delta p_{n}/p_{n}$ in comparison to the exact numerical results, which shows that the discrepancy for low levels is mitigated by high-order perturbative calculations.
 In the vicinity of the bottom, we can apply the harmonic approximation ($\xi_{kn}=0$). Under this approximation, the ratio of probabilities over adjacent levels is
@@
 The zero-order term yields a constant effective temperature.
-When we include the correction to the order of $\lambda$, the correction leads to changes in $N_{eff}$ showing a good agreement with the numerical results for levels near the bottom.
+When we include the correction up to the order of $\lambda$, the correction leads to changes in $N_{eff}$ showing a good agreement with the numerical results for levels near the bottom.
 Figure 7: Relationship between effective frequency $\omega_{eff}=\Delta E_{1}$ and damping strength $\kappa$.
@@
 Figure 8: Renormalized effective Bose distribution $\tilde{N}$ due to dephasing effects for (a) the LAS and (b) the HAS.
-We compare the theoretical predictions from Eq.
+We compare theoretical predictions from Eq.
 (36) with the exact numerical results and extract the linear relationships.
 III.2 Strong Damping and Dephasing
-In this section, we explore the dynamics and stationary state of the system under conditions of strong damping and dephasing, which can significantly alter the behavior predicted by the weak damping approximation.
+In this section, we explore the dynamics and the stationary state of the system under conditions of strong damping and dephasing, which can significantly alter the behavior predicted by the weak damping approximation.
 III.2.1 Strong damping
@@
 Instead, the level spacing undergoes slight renormalization for strong damping.
-Such effect can be observed through the emission spectrum $S(\omega)$, representing the spectral density of photons emitted by the driven resonator, is given by
+Such an effect can be observed through the emission spectrum $S(\omega)$, which represents the spectral density of photons emitted by the driven resonator and is given by
 $S(\omega)=2\operatorname{Re}\int_{0}^{\infty}e^{-i\omega t}\text{Tr}[a^{\dagger}e^{\mathcal{L}t}(a\rho_{\text{st}})]{\rm d}t\,.$ (34)
@@
 (29).
-Substituting them into Eq.
+Substituting these parameters into Eq.
 (26), we obtain the effective frequency $\omega_{eff}=\Delta E_{1}$.
@@
 (34) with those from our perturbation theory, demonstrating satisfactory consistency.
-The minor discrepancy arises primarily from higher-level spacings $\Delta E_{n}(n>1)$, which are typically smaller than the first-level spacing $\Delta E_{1}$.
+The minor discrepancy arises primarily from higher-level spacings $\Delta E_{n}(n>1)$, which are generally smaller than the first-level spacing $\Delta E_{1}$.
 For more accurate results, the average of all level spacings should be considered in the calculations.
@@
 For convenience, we define a generalized Lindblad operator ${\cal L}[A;B]\rho\equiv 2A\rho B-BA\rho-\rho AB$.
-In the spirit of the rotating wave approximation, we obtain the renormalized master equation for the displaced and squeezed density ioperator $\bar{\rho}=S^{\dagger}D^{\dagger}{\rho}DS$ (see detailed derivation in Appendix B)
+In the spirit of the rotating wave approximation, we obtain the renormalized master equation for the displaced and squeezed density operator $\bar{\rho}=S^{\dagger}D^{\dagger}{\rho}DS$ (see the detailed derivation in Appendix B)
 $\frac{d\bar{\rho}}{dt}$ $=$ $-i[\bar{H},\bar{\rho}]+\frac{\kappa}{2}\{(1+\tilde{N}){\cal D}[a]\bar{\rho}+\tilde{N}{\cal D}[a^{\dagger}]\bar{\rho}\}$ (35)
@@
 $=$ $|u|^{2}+\bar{n}(|u|^{2}+|v|^{2})+\frac{\eta^{ph}}{\kappa}|\alpha^{*}u+\alpha v^{*}|^{2}\,.$
-We verify our predictions for renormalized Bose distribution by comparing them with exact numerical simulations according to the probabilities over the two lowest levels, specifically $\tilde{N}=p_{2}/(p_{1}-p_{2})$.
+We verify our predictions for the renormalized Bose distribution by comparing them with exact numerical simulations according to the probabilities over the two lowest levels, specifically, $\tilde{N}=p_{2}/(p_{1}-p_{2})$.
 In Fig.
-8, we plot and extract the renormalized Bose distribution as a function of the dephasing, which demonstrates an excellent agreement between the numerical results and the analytical predictions given by Eq.
+8, we plot and extract the renormalized Bose distribution as a function of dephasing, which demonstrates an excellent agreement between the numerical results and the analytical predictions given by Eq.
 (36) for both LAS and HAS.
@@
 In this work, we have investigated the quantum dissipative dynamics of a driven Duffing oscillator near the bottoms of its stable states.
-We elucidated the intricate interplay between the nonlinearity, quantum fluctuations, and the influence of an external driving field.
+We elucidated the intricate interplay among the nonlinearity, quantum fluctuations, and the influence of an external driving field.
 We formulated an effective quantum master equation that encompasses quantum and thermal fluctuations, strong damping, and dephasing within a unified framework.
-We have developed a refined perturbation theory to analyze the quantum dynamics near both the LAS and the HAS of the Duffing oscillator.
-While the LAS behavior can be approximated using a harmonic oscillator model, the HAS exhibits more complex behavior due to the significant nonlinear terms.
-Due to quantum fluctuations, even at zero temperature, higher energy levels near the bottom of the potential well are excited.
-We calculated the level spacing and effective temperature near the bottom of the HAS and compare them with numerical simulations, demonstrating the accuracy and utility of our proposed approach.
+We have developed a refined perturbation approach to analyze the quantum dynamics near both the LAS and the HAS of the Duffing oscillator.
+While the LAS behavior can be approximated using a harmonic oscillator model, the HAS exhibits more complex behavior owing to significant nonlinear terms.
+Because of quantum fluctuations, even at zero temperature, higher energy levels near the bottom of the potential well are excited.
+We calculated the level spacing and effective temperature near the bottom of the HAS and compared them with numerical simulations, demonstrating the accuracy and utility of our proposed approach.
 We also investigated the effects of strong damping and dephasing on the system’s dynamics.
@@
 ACKNOWLEDGMENTS
-This work was supported by Natural Science Foundation of China (Grant No. 12475025).
+This work was supported by the Natural Science Foundation of China (Grant No. 12475025).
 AUTHOR DECLARATIONS
@@
 $=\gamma^{2}(\sum_{l\neq n}\epsilon^{(0)}_{n}\xi^{(2)}_{ln}|\psi^{(0)}_{l}\rangle+\sum_{k\neq n}\xi^{(1)}_{kn}\epsilon^{(1)}_{n}|\psi^{(0)}_{k}\rangle+\epsilon^{(2)}_{n}|\psi^{(0)}_{n}\rangle),$
-which gives us the second order perturbative result:
+which gives us the second-order perturbative result:
 $\epsilon^{(2)}_{n}$ $=$ $\sum_{k\neq n}\langle\psi^{(0)}_{n}|H_{1}|\psi^{(0)}_{k}\rangle\xi^{(1)}_{kn}+\langle\psi^{(0)}_{n}|H_{2}|\psi^{(0)}_{n}\rangle,$